OpenStudy (aye_my_peeps):

a gardener weeds 2/3 of a flower bed in 6 mins. how much of the flower bed did the gardener weed in 1 min?

OpenStudy (aye_my_peeps):

@truwhovian11

OpenStudy (minecraftqueen):

There are multiple ways you can solve this: (2/3)/6=2/18=1/9 or 2/3 in 6 minutes I find how long it takes to finish a third which is 6/2=3 minutes I multiply 3 min by 3 to find how long it takes to finish 3/3 of the garden This tells me that 1 garden=9 minutes How much garden is done in 1 minute is equal to 1/9